OpenStudy (anonymous):

A balloon is filled with helium at sea level. The volume was measured to be 3.1 liters at a pressure of 0.97 atmosphere. If the balloon is released, it drifts upward. At a height of 4 kilometers, the pressure is measured to be 0.61 atmospheres. What is the volume of the balloon

OpenStudy (anonymous):

I think you can apply boyle's law which states that product of pressure and volume is a constant. p1*v1=p2*v2 p1=.97atm v1=3.1L p2=.61 v2=? do the math.